# Cross-Modal Geometric Hierarchy Fusion: An Implicit-Submap Driven Framework for Resilient 3D Place Recognition

### v1

@@ IV-C Experiment Performance Comparisons
To comprehensively test our method, we evaluated its performance in two major scenarios: short-term and long-term relocalization. Here, ”short-term” refers to real-time detection of loop closure points, while ”long-term” denotes the ability to recognize locations when revisiting the same area after a long period, specifically more than one day.
  1. 1)
@@
    short-term: The PR curves for our proposed method and existing methods are shown in Fig. 5. We compared a total of 20 trajectories. Our method demonstrates superior performance compared to the existing methods. In the KITTI00 sequence, since there are only small rotational and translational changes when passing through the same location, most methods perform well. However, in scenarios with significant rotational and translational changes at the same location – such as KITTI05, KITTI08, KITTI360_04, and KITTI360_05 – existing methods tend to generate inaccuracies, whereas our method retains robust performance. When dealing with numerous similar/repetitive scenes or revisiting the same location multiple times (e.g., the six sequences in the MulRan dataset), all methods show some susceptibility to these challenging conditions. Nevertheless, our method still maintains good performance. Table II presents the AUC values and peak F1 scores. Our method surpasses the compared approaches. To provide a detailed analysis of each method, two sets of sequences for testing are presented in the figure. All methods display their optimal performance (indicated by the maximum F1 score). Compared to other methods, ours achieves higher True Positives (TP) and lower False Positives (FP) and False Negatives (FN). To further highlight the distinction, we selected KITTI02 and MulRan04 tracks and visualized their place recognition results (including TP, FP, FN, and TN) based on each method’s optimal F1 score threshold. As illustrated in Figs. 6 and 7, it is evident that our approach generates fewer false matches and better detects loopback points, thereby delivering superior performance.
  2. 2)
@@
    long-term: A robust place recognition system should perform reliably despite long-term environmental variations. To evaluate long-term place recognition performance, multiple intermittent datasets covering the same area were selected as the map sequence and query sequence, respectively. The results of our recall-precision experiments are shown in Fig. 8. AUC and F1 scores are presented in Table III. Our method demonstrates superior performance across all metrics. Although performance declines compared to short-term experiments on the same dataset, this results from the expanded retrieval scope and environmental changes inherent to long-term tasks, which increase place recognition complexity.
As demonstrated, our method outperforms comparative approaches. The key factors contributing to this superior performance are threefold. First, our implicit representation ensures enhanced uniformity in the acquired BEV, which directly translates to more consistent spatial descriptors. Second, our geometric descriptors strategically disregard ground-level information (prone to redundancy) while focusing exclusively on regions containing rich geometric features such as edges and corners—a distinct methodological departure. Finally, the implicit representation inherently suppresses dynamic elements, enabling effective noise interference elimination.
TABLE III: AUC AND MAX F1 SCORES ON TESTING SEQUENCES.
Method NCLT-L1 NCLT-L2 MulRan-L1 MulRan-L2
@@
TABLE IV: AUC AND MAX F1 SCORES ON ABLATION STUDY.
Implicit Representation BEV Image Normal Vector KITTI KITTI-360 NCLT Mulran
0.91/0.86 0.85/0.84 0.8/0.79 0.51/0.60
0.59/0.57 0.50/0.48 0.47/0.42 0.32/0.27
0.76/0.79 0.75/0.77 0.73/0.71 0.42/0.40
0.81/0.79 0.72/0.73 0.76/0.78 0.48/0.50
Note: On the left is the AUC and on the right is the max F1 score. The best-performing method is bolded. The individual datasets are presented as their average levels.
@@ IV-D Ablation Study
To validate the contribution of the point-based implicit representation to our place recognition framework, we conducted two experimental configurations on our dataset. The first configuration retained all default parameters, while the second removed the implicit representation module. In the absence of an implicit submap, we adopted a voxelization-based statistical method to generate occupancy grids for standard submaps. Specifically, we calculated point density within each voxel and applied a density threshold matching that of implicit submaps to ensure resolution consistency. For normal vector estimation, we employed a principal component analysis (PCA)-based approach: PCA was computed on each point’s local neighborhood, with the eigenvector corresponding to the smallest eigenvalue designated as the surface normal. Ablation tests were performed across four datasets covering short-term and long-term scenarios. Quantitative results (AUC and Maximum F1-score) for all datasets are summarized in Table IV.
Building on the implicit submap representation, we conducted two additional experiments: one utilizing exclusively BEV features and the other using solely normal vector information features. Detailed quantitative comparisons are provided in Table V. As evidenced in the table, removing the implicit submap representation causes marked performance degradation. This decline stems primarily from two sources: 1) errors in surface normal estimation, and 2) inaccuracies in occupancy grid generation, both of which introduce noise that compromises 3D place recognition robustness. It can also be found that only the macro-level information from the BEV perspective and the local micro-scale information on the main 3D segments cooperate to produce a robust descriptor.
TABLE V: TIME CONSUMPTION (MS) OF DIFFERENT METHOD.
@@
IV-E Computational Cost
We conducted comparative experiments on the runtime of our method to demonstrate its efficiency in 3D place recognition. We statistically analyzed the average time required to generate descriptors for each submap across all sequences, including calculating the average time for the three most critical components, and compared our method with other approaches. The running time of our method was calculated by measuring the time for descriptor generation and retrieval. The results are presented in Table V. From the table, we can see that our method is not worse than most methods in running time, even if we use implicit expressions. Our implicit representation is lightweight and efficient. We add the Voxel Hashing technique to the process of obtaining nerve points to index nerve points efficiently and reduce some unnecessary optimization iterations. Meanwhile, our method for segmenting major 3D segments is also fast and efficient. Finally, real-time performance is achieved.
@@ V CONCLUSION
In this paper, we propose a novel 3D place recognition method inspired by implicit representations. Leveraging submap-based implicit representations, the method begins by acquiring only occupancy grids, uniform triangular points, and corresponding globally consistent normal vectors. This information is then collaboratively utilized to generate BEV images and information on clustered 3D segments, depending on exact meaning. We employ Log-Gabor filters to derive descriptors for the BEV images and compute descriptors for the 3D segments based on the angular differences between their normal vectors. By concatenating these two descriptors, we construct a cross-modal geometric hierarchy fusion feature. To the best of our knowledge, this is the first framework that utilizes implicit representations to fuse cross-modal geometric features for 3D place recognition, significantly advancing the exploitation of implicit representations’ potential in 3D place recognition. In future work, we will further explore the possibilities and roles of implicit representations in place recognition, including the development of new representation paradigms specifically tailored for this task.

### v3

@@ IV-C Experiment Performance Comparisons
To comprehensively test our method, we evaluated its performance in two major scenarios: short-term and long-term relocalization. Here, ”short-term” refers to real-time detection of loop closure points. ”Long-term” is the ability to recognize locations when revisiting the same area after more than one day. We further benchmark our approach against several learning-based counterparts. Our method ingests sub-maps accumulated from multi-frame point clouds. In contrast, most learning-based techniques operate on single-frame inputs. Therefore, we conduct the comparison under explicitly aligned data regimes to ensure fairness. For a rigorous comparison, we re-formulate the learning-based baselines (LCDnet[84], Bevplace++[85]) so they ingest multi-frame inputs generated under the same conditions as our method. We adjust the temporal window and voxelization parameters so the tensors are shape-compatible with the pre-trained weights. This enables direct evaluation without retraining.
@@
1) short-term: The PR curves for our proposed method and existing methods are shown in Fig. 5. We compared a total of 20 trajectories. Our method demonstrates superior performance compared to the existing methods. In the KITTI00 sequence, there are only small rotational and translational changes when passing through the same location, so most methods perform well. In contrast, scenarios with significant rotational and translational changes at the same location (such as KITTI05, KITTI08, KITTI360-04, and KITTI360-05) cause existing methods to generate inaccuracies. Our method retains robust performance in these cases. When there are numerous similar or repetitive scenes or when revisiting the same location multiple times, as in the six sequences in the MulRan dataset, all methods show some susceptibility to these challenging conditions. Nevertheless, our method still maintains good performance. Table II presents the AUC values and peak F1 scores. Our method surpasses the compared approaches. Two sets of sequences for testing are presented in the figure to provide a detailed analysis of each method. All methods display their optimal performance, which is indicated by the maximum F1 score. Compared to other methods, ours achieves higher True Positives (TP) and lower False Positives (FP) and False Negatives (FN). To further highlight the distinction, we selected KITTI02 and MulRan04 tracks. We visualized their place recognition results, including TP, FP, FN, and TN, based on each method’s optimal F1 score threshold. As shown in Figs. 6 and 7 show that our approach generates fewer false matches and detects loopback points more effectively. This delivers superior performance.
@@
2) long-term: A robust place recognition system should perform reliably despite long-term environmental variations. To evaluate long-term place recognition performance, multiple intermittent datasets covering the same area were selected as the map sequence and query sequence, respectively. The results of our recall-precision experiments are shown in Fig. 8. AUC and F1 scores are presented in Table III. Our method demonstrates superior performance across all metrics. Although performance declines compared to short-term experiments on the same dataset, this decline is largely attributed to increased retrieval scope and significant environmental changes over time, such as seasonal differences, lighting shifts, or structural modifications. These factors make the long-term place recognition task more complex.
3) learning-based: Although our framework leverages a learned implicit representation, the descriptor itself is constructed in a fully heuristic manner. This obviates the need for place-recognition labels and eliminates any training overhead. Nonetheless, to situate our approach within the state of the art, we benchmark it against leading learning-based baselines. The resulting PR curves are reported in Fig. 9. LCDNet releases weights exclusively for KITTI-360, and BecPlace++ provides a reproducible training recipe on KITTI. We benchmark our method against BecPlace++ on KITTI-02 and against LCDNet on KITTI-360-00, thereby adhering to the publicly available model domains. As shown, our method surpasses LCDNet on KITTI-360-00 while remaining marginally behind BEVPlace++ on KITTI-02, particularly at higher recall regions. This clarifies the contexts in which our approach’s performance is slightly lower, corroborating both its strong empirical efficacy and its untapped potential.
As demonstrated, our method surpasses comparative approaches for four reasons. First, our implicit representation enhances BEV uniformity, producing more consistent spatial descriptors. Second, we integrate information from multiple modalities at both macro and micro scales, then fuse them geometrically, resulting in a richer and more hierarchical understanding of the environment. Third, our geometric descriptors disregard redundant ground-level data and focus on distinctive features, such as edges and corners, marking a clear methodological difference. Finally, the implicit representation suppresses dynamic elements, effectively reducing noise interference.
TABLE III: Performance Comparison on Long-Term Sequences (AUC $\uparrow$ / Max F1 Score $\uparrow$ )
Method NCLT MulRan
L1 L2 L1 L2
@@
TABLE IV: Ablation Study Results: AUC $\uparrow$ and Max F1 $\uparrow$ Scores across Datasets
Components KITTI KITTI-360 NCLT Mulran
Implicit Rep. BEV Image Normal Vector (AUC / F1) (AUC / F1) (AUC / F1) (AUC / F1)
0.91 / 0.86 0.85 / 0.84 0.80 / 0.79 0.51 / 0.60
0.59 / 0.57 0.50 / 0.48 0.47 / 0.42 0.32 / 0.27
0.76 / 0.79 0.75 / 0.77 0.73 / 0.71 0.42 / 0.40
0.81 / 0.79 0.72 / 0.73 0.76 / 0.78 0.48 / 0.50
Note: The best performance for each dataset is highlighted in bold. Results are presented as AUC score / Max F1 score.
Figure 9: Precision–recall curve KITTI02, KITTI360-00 on learning-based method.
@@ IV-D Ablation Study
To validate the contribution of the point-based implicit representation to our place recognition framework, we conducted two experimental configurations on our dataset. The first configuration retained all default parameters. The second configuration removed the implicit representation module. Without an implicit submap, we used a voxelization-based statistical method to generate occupancy grids for standard submaps. We calculated the point density within each voxel and set a density threshold that matched that of the implicit submaps. This ensured resolution consistency. For normal vector estimation, we used a principal component analysis (PCA)-based approach. PCA was computed on each point’s local neighborhood. The eigenvector with the smallest eigenvalue was designated as the surface normal. Ablation tests were conducted across four datasets, encompassing both short-term and long-term scenarios. Quantitative results (AUC and Maximum F1-score) for all datasets are summarized in Table IV.
Building on the implicit submap representation, we conducted two additional experiments. One used only BEV features, while the other used only normal vector features. Detailed quantitative comparisons are provided in Table IV. As shown in the table, removing the implicit submap representation leads to significant performance degradation. This decline is due mainly to two sources: errors in surface normal estimation and inaccuracies in occupancy grid generation. Both introduce noise that compromises 3D place recognition robustness. The results also show that macro-level BEV information and local micro-scale information from the main 3D segments work together to produce a robust descriptor.
TABLE V: TIME CONSUMPTION (MS) OF DIFFERENT METHODS
@@
TABLE VI: Memory consumption in MB
Method Raw point cloud Ours
KITTI02 9002.3 90.1(1.0%)
KITTI06 2052.2 18.58(0.9%)
KITTI360-00 21210.3 255.3(1.1%)
NCLT02 15366 108.2(0.07%)
IV-E Computational Cost and Storage Efficiency
We conducted comparative experiments on the runtime of our method to demonstrate its efficiency in 3D place recognition. We statistically analyzed the average time required to generate descriptors for each submap across all sequences. For the three most critical components, we calculated their average times separately and compared our method with other approaches. The running time of our method was calculated by measuring the times required for descriptor generation and retrieval. The results are presented in Table V. From the table, we can see that our method is not inferior to most methods in terms of running time, even when using implicit expressions. Our implicit representation is lightweight and efficient. We use the Voxel Hashing technique to efficiently index nerve points and reduce unnecessary optimization iterations. Our segmentation of the major 3D segments is also fast and efficient. Finally, our method achieves real-time performance.
We further demonstrate the storage efficiency advantage conferred by our implicit neural-point representation factor. Prior work often focuses solely on descriptor extraction while retaining raw point-cloud formats for downstream map fusion and maintenance. On several representative sequences, we quantify the compression ratio between our implicit encoding and the raw point clouds. Table VI shows that our representation yields substantial storage savings. This advantage unlocks significant potential for subsequent map-merging and edge-cloud collaborative mapping tasks.
@@ V CONCLUSION
In this paper, we propose a novel 3D place recognition method inspired by implicit representations. The method begins by acquiring only occupancy grids, uniform triangular points, and globally consistent normal vectors. Submap-based implicit representations are leveraged to process this information. The data is then used to generate BEV images and information on clustered 3D segments. We employ Log-Gabor filters to derive descriptors for the BEV images. We also compute descriptors for the 3D segments using the angular differences between their normal vectors. By concatenating these two descriptors, we construct a cross-modal geometric hierarchy fusion feature. To our knowledge, this is the first framework to introduce cross-modal geometric features for 3D place recognition. It leverages the expressive power of implicit representations to establish a new performance frontier.
In future work, we will further explore the possibilities and roles of implicit representations in place recognition. This includes developing new representation paradigms specifically tailored for this task. Additionally, we note that the current method has not been evaluated under a unified dataset benchmark. We aim to construct a large-scale, standardized dataset that unifies and facilitates fair comparisons across different environments. Based on this unified benchmark, we will also attempt to design an end-to-end place recognition framework fully based on implicit representations.